OpenStudy (nykoreyan):

What is the solution of the system of equations? y = –4x + 10 y = –2x – 6 (8, –22) (–22, 8) (–2.67, 20.67) (–2, –2)

OpenStudy (nykoreyan):

@JasonSpade

OpenStudy (denonakavro):

It is 8,-22

OpenStudy (anonymous):

A.(8,-22)

OpenStudy (er.mohd.amir):

y=y -4x+10=-2x-6 2x=16 x=8 y=-2*8-6=-22